In a staff model HMO (Health Maintenance Organization), the requirement for MDs' (Medical Doctor) time is estimated to be 185 hours per day. The utilization target for MDs is 90 percent. Human Resources' benefit plan for them includes eleven holidays, ten sick days,and twenty-one paid vacation days per year.

Determine the core level of MD FTEs (Full-time equivalent) for the HMO.
Answer #1

Measure of work in one year = 11000 x 365 = 4015000 mins

considering 90% usage target, in the event that 90% = 4015000m mins, at that point.

100% = (4015000x100)/90

= 446111.11 min

no. of day representatives will work considering occasions days off and paid get-away = (365-11-10-21) = 323 days

presently 446111.11 mins work to be done in 323 days

=13822.489 mins ordinary

= 13822.489/60 = 230.191 hours daily

considering FTE's will labor for 24 hours than 230.191/24 = 10 MD FTE's for the HMO